4. Suppose that the inverse demand function for movies is p = 120 – Q1 for college students and p = 120 – 2Q2 for other town residents. What is the town’s total demand function (Q = Q1 + Q2 as a function of p)?


1
Expert's answer
2021-01-19T07:22:43-0500

The total demand function for the town is the sum of the demand functions for the two groups.

Let suppose that "P=100". Then, from the first and second demand functions we get:


"Q_1=120-100=20,""Q_2=\dfrac{120-100}{2}=10."

Therefore, "Q=Q_1+Q_2=20+10=30."

Then, let's do the same calculations for "P=60":


"Q_1=120-60=60,""Q_2=\dfrac{120-60}{2}=30,""Q=Q_1+Q_2=60+30=90."

As we can see from calculations, we obtain two points on total demand curve: ("x_1", "y_1") and ("x_2", "y_2") or (30, 100) and (90, 60), respectively. Let's use the point slope formula to find the total demand function:


"y-y_1=m(x-x_1),"

here, "y=P", "y_1=100", "x=Q", "x_1=30", "m" is the slope of the line.

So, we can rewrite the equation:


"(P-100)=m(Q-30)."

The slope of the line can be found as follows:


"m=\dfrac{\Delta y}{\Delta x}=\dfrac{60-100}{90-30}=-\dfrac{40}{60}."

Then, substituting the slope into the equation of the line, we get:


"(P-100)=-\dfrac{40}{60}(Q-30),""3P-300=60-2Q,""P=120-0.67Q."

Answer:

"P=120-0.67Q."